OpenStudy (anonymous):

Let \[ a_{0}=1,\; a_{1}=2,\; a_{2}=24,\; a_{n}=\frac{ 6a_{n-1}^{2}a_{n-3}-8a_{n-1}a_{n-2}^{2}}{a_{n-2}a_{n-3}}. \] Show that every element of this sequence is an integer.

OpenStudy (asnaseer):

\[a_{n}=\frac{ 6a_{n-1}^{2}a_{n-3}-8a_{n-1}a_{n-2}^{2}}{a_{n-2}a_{n-3}}\]\[\qquad=2a_{n-1}(\frac{ 3a_{n-1}a_{n-3}-4a_{n-2}^{2}}{a_{n-2}a_{n-3}})\]\[\qquad=2a_{n-1}(\frac{3a_{n-1}}{a_{n-2}}-\frac{4a_{n-2}}{a_{n-3}})\]therefore:\[\frac{a_n}{a_{n-1}}=2(\frac{3a_{n-1}}{a_{n-2}}-\frac{4a_{n-2}}{a_{n-3}})\]now let:\[r_n=\frac{a_n}{a_{n-1}}\]therefore:\[r_n=2(3r_{n-1}-4r_{n-2})\]and I'm sure we can show that all \(r_n\) are also integers using induction as we know from the given values that:\[r_1=\frac{2}{1}=2\]\[r_2=\frac{24}{2}=12\]this would then prove that all \(a_n\) are also integers (I think).